# Serializing Java Objects in Plain Code

### v3

@@ 4.3 RQ3 (Performance)
{mdframed}
[nobreak=true,style=mpdframe,frametitle=Answer to RQ3] The performance evaluation of ProDJ shows that plain-code serialization takes a few milliseconds per serialized object, which is invisible to the user in interactive use-cases. Contrary to traditional serialization techniques, plain-code serialization involves a computation cost before and after execution.
@@ 4.4 RQ4 (User Study)
We conduct a user study per the protocol defined in subsection 3.4, based on an online survey with developers. The goal of the survey is to assess the opinion of developers regarding the readability of tests containing plain-code serialized objects. This survey received 17 responses by developers with 20 years of median experience with programming, and 13 years programming with Java. We prepare 16 pairs of tests in total, presenting five random pairs to each respondent. We select two tests for each project which have less than 30 lines of Java code, including outlined utility methods. For the JSON and XML variants of these tests, we collapsed identical lines in the JSON and XML code. The maximum number of lines in the tests with JSON and XML tests was 1,797 and 1,787, respectively. Each respondent was shown five randomly selected pairs. One test within the pair contains plain-code serialized objects, while the other was the same test, but with either JSON- or XML-serialized objects. We randomized the order in which the plain-code and JSON or XML tests were presented. The respondents were requested to mark which test in each pair they preferred, with respect to its readability. Moreover, 11 of the 17 respondents shared additional qualitative comments about serialization.
Project Test pair Votes C or F kappa (agreement)
Plain-code JSON/XML
PDFBox PDF1–JSON 8 0 F 1.0 (perfect)
PDF1–XML 3 0 F 1.0 (perfect)
PDF2–JSON 4 2 F -0.07 (poor)
PDF2–XML 6 1 F 0.43 (moderate)
GraphHopper GH1–JSON 7 1 F 0.5 (moderate)
GH1–XML 6 1 F 0.43 (moderate)
GH2–JSON 2 0 C 1.0 (perfect)
GH2–XML 6 0 F 1.0 (perfect)
BroadLeaf BL1–JSON 4 1 F 0.2 (slight)
BL1–XML 9 0 F 1.0 (perfect)
BL2–JSON 5 0 F 1.0 (perfect)
BL2–XML 2 0 C 1.0 (perfect)
Gephi GEP1–JSON 4 0 F 1.0 (perfect)
GEP1–XML 6 0 F 1.0 (perfect)
GEP2–JSON 3 0 F 1.0 (perfect)
GEP2–XML 4 0 F 1.0 (perfect)
Total 16 79 6
@@
Table 5: Summary of our user study. We select 2 tests from each of the four projects for the study. The second column lists the 16 test pairs we prepare using these tests. Each pair contains one test with plain-code serialized objects and an equivalent test with either JSON or XML representations of the same serialized objects. We present 5 random test pairs to each of the 17 participants of the study, who vote for tests with plain-code (column 3) or JSON/XML (column 4) serialized objects, based on their readability. The last column presents the Cohen’s (C) or Fleiss’ (F) kappas of agreement between the participants.
@@
Table 5 summarizes the responses from the survey. We find that through 79 of the total 85 votes (92.9%), the developers showed a preference for the tests containing the plain-code serialized objects. In only 6 cases, they preferred the versions with XML or JSON representations. For 11 of the 16 test pairs, the preference for plain-code serialized objects was unanimous. In those cases, the agreement measure (Cohen’s and Fleiss’ kappas) is a perfect 1.0. In the 5 remaining cases, the agreement varies from poor to moderate.
@@
[…]
[4pt]       I [prefered] the shorter versions because the longer ones contained serialization ”noise” that did not seem relevant.
[…]
[4pt]       Intuitively, the [plain-code version] is often more readable. However, legacy code may be hard to test as classes could have many dependencies that are hard to construct. In those cases, the readability gap between object construction and deserialization could be narrowed.
@@
[…]
[4pt]       I prefer the unchanging nature of JSON and XML, which do not get refactored automatically in sync with the application code, when the classes under test change.
@@
[…]
[4pt]       I believe I would store the long text-snippets as variables in another file, instead of having them in the test.
@@
[…]
[4pt]       When sending data back and forth over the wire, more efficient formats may be better, like Protobuf […]. Overall, it really depends on the task at hand.
This is valid point Our technique is entirely optimized for readability, and we do recommend binary serialization for performance-sensitive use cases.
{mdframed}
[nobreak=true,style=mpdframe,frametitle=Answer to RQ4] Our user study demonstrates, through a sound protocol, that developers rate tests containing plain-code serialized objects higher with respect to their readability, in comparison with equivalent tests with XML- or JSON-serialized objects.

### v4

@@ 4.3 RQ3 (Performance)
The performance evaluation of ProDJ shows that plain-code serialization takes a few milliseconds per serialized object, which is invisible to the user in interactive use-cases. Contrary to traditional serialization techniques, plain-code serialization involves a computation cost before and after execution.
@@ 4.4 RQ4 (User Study)
We conduct a user study per the protocol defined in subsection 3.4, based on an online survey with developers. The goal of the survey is to assess the opinion of developers regarding the readability of tests containing plain-code serialized objects. This survey received 17 responses from developers who have between 8 and 43 years of experience with programming (median 20), and between 1 and 30 years programming with Java (median 13). We prepare 16 pairs of tests in total, presenting five random pairs to each respondent. We select two tests for each project which have less than 30 lines of Java code, including outlined utility methods. From Table 5, we see that one test each in PDFBox (PDF1) and GraphHopper (GH1) have 23 lines of code (LOC), the maximum among the 8 plain-code tests generated by ProDJ. For the JSON and XML variants of these tests, we collapsed identical lines in the JSON and XML code. The maximum number of lines in the tests with JSON and XML tests is 1,794 and 1,784, respectively (both for PDF2 in Table 5). Note from the table that in each of the 16 test pairs, as well as across all the pairs, the number of lines of code for the plain-code test is significantly lesser than its corresponding JSON and XML representation. Each respondent was shown five randomly selected pairs. One test within the pair contains plain-code serialized objects, while the other was the same test, but with either JSON- or XML-serialized objects. We randomized the order in which the plain-code and JSON or XML tests were presented. The respondents were requested to mark which test in each pair they preferred, with respect to its readability. Moreover, 11 of the 17 respondents shared additional qualitative comments about serialization.
Project Test pair LOC Votes C or F kappa (agreement)
Plain-code JSON XML Plain-code JSON/XML
PDFBox PDF1–JSON 23 76 - 8 0 F 1.0 (perfect)
PDF1–XML - 65 3 0 F 1.0 (perfect)
PDF2–JSON 8 1,794 - 4 2 F -0.07 (poor)
PDF2–XML - 1,784 6 1 F 0.43 (moderate)
GraphHopper GH1–JSON 23 349 - 7 1 F 0.5 (moderate)
GH1–XML - 259 6 1 F 0.43 (moderate)
GH2–JSON 18 192 - 2 0 C 1.0 (perfect)
GH2–XML - 147 6 0 F 1.0 (perfect)
BroadLeaf BL1–JSON 8 24 - 4 1 F 0.2 (slight)
BL1–XML - 22 9 0 F 1.0 (perfect)
BL2–JSON 10 20 - 5 0 F 1.0 (perfect)
BL2–XML - 17 2 0 C 1.0 (perfect)
Gephi GEP1–JSON 11 1,128 - 4 0 F 1.0 (perfect)
GEP1–XML - 704 6 0 F 1.0 (perfect)
GEP2–JSON 11 1,128 - 3 0 F 1.0 (perfect)
GEP2–XML - 704 4 0 F 1.0 (perfect)
Total 16 112 4,711 3,702 79 6
@@
Table 5: Summary of our user study. We select 2 tests from each of the four projects for the study. The second column lists the 16 test pairs we prepare using these tests. Each pair contains one test with plain-code serialized objects and an equivalent test with either JSON or XML representations of the same serialized objects. The number of lines of code in each test in the pair is indicated in the column LOC. We present 5 random test pairs to each of the 17 participants of the study, who vote for tests with plain-code (column 3) or JSON/XML (column 4) serialized objects, based on their readability. The last column presents the Cohen’s (C) or Fleiss’ (F) kappas of agreement between the participants.
@@
Table 5 summarizes the responses from the survey. We find that through 79 of the total 85 votes (92.9%), the developers showed a preference for the tests containing the plain-code serialized objects, regardless of the differences in the number of lines of code between the plain-code and JSON/XML variants. In only 6 cases, developers preferred the versions with XML or JSON representations. For 11 of the 16 test pairs, the preference for plain-code serialized objects was unanimous. In those cases, the agreement measure (Cohen’s and Fleiss’ kappas) is a perfect 1.0. In the 5 remaining cases, the agreement varies from poor to moderate.
@@
                                           I [prefered] the shorter versions because the longer ones contained serialization ”noise” that did not seem relevant.
                                           Intuitively, the [plain-code version] is often more readable. However, legacy code may be hard to test as classes could have many dependencies that are hard to construct. In those cases, the readability gap between object construction and deserialization could be narrowed.
@@
                                           I prefer the unchanging nature of JSON and XML, which do not get refactored automatically in sync with the application code, when the classes under test change.
@@
                                           I believe I would store the long text-snippets as variables in another file, instead of having them in the test.
@@
                                           When sending data back and forth over the wire, more efficient formats may be better, like Protobuf […]. Overall, it really depends on the task at hand.
This is valid point. Our technique is entirely optimized for readability, and we do recommend binary serialization for performance-sensitive use cases.
Our user study demonstrates, through a sound protocol, that developers rate tests containing plain-code serialized objects higher with respect to their readability, in comparison with equivalent tests with XML- or JSON-serialized objects.
@@ 5 Limitations and Threats
A threat to the validity of our conclusions from the user study of RQ4 arises from the expectation that developers may automatically prefer shorter versions of tests with plain-code serialized objects over their XML or JSON counterparts. We design our user study to minimize this threat and also report the deviations we observe from this outcome. Our key finding is that, even for test pairs with comparable number of lines of code, developers prefer plain-code serialized objects over JSON or XML.